# From Tabula Rasa to Emergent Abilities: Discovering Robot Skills via Real-World Unsupervised Quality-Diversity

### v1

@@ 1 Introduction
In this work, we introduce Unsupervised Real-world Skill Acquisition (URSA), a framework for unsupervised QD that operates directly on physical robots. URSA extends the Quality-Diversity Actor-Critic (QDAC) framework for unsupervised learning and leverages the DayDreamer [13] world model as its backbone, enabling efficient skill acquisition directly on hardware without relying on simulation or sim-to-real transfer. We evaluate URSA on a Unitree A1 quadruped and show it outperforms baselines in 5 out of 9 simulated and 3 out of 5 real-world damage scenarios. This demonstrates its potential for enabling adaptable, resilient skill discovery directly on physical robots. Our key contributions are:
@@ 4.2 Efficient Skill Sampling from the Reachable Space
To maximize the diversity of achieved skills, URSA must efficiently explore the reachable skill space $\mathcal{Z}_{p}$ while ensuring safe execution. URSA addresses this by maintaining a diversity-aware repertoire $\mathcal{R}$ and sampling skills from an flexible surrogate distribution $q(\cdot)$ learned from experience. Figure 2 and Algorithm 1 provide high-level overviews of the skill collection and sampling process.
@@
As the repertoire expands, the typical distance between skills increases, requiring adaptive tuning of the constraint threshold $\delta$ in Problem P3. To that end, we introduce a hyperparameter, which controls the number of distinct skills that the robot can excute, $N_{\boldsymbol{z}}$. We compute $\delta$ as the mean nearest-neighbor distance, if the repertoire contained $N_{\boldsymbol{z}}$ uniformly distributed skills (see equation in Appendix B.5). This ensures that executing one skill does not inadvertently satisfy multiple constraints.
@@ 5.1 Experimental Setup and Baselines
We evaluate URSA on the Unitree A1 quadruped in both PyBullet simulation and real-world environments. The robot state includes joint angles and velocities for its 12 joints; actions specify target joint positions executed via a 20 Hz PD controller. Skills are sampled every 250 steps from a fixed repertoire of $N_{R}=4096$. Training in simulation runs for 1 million timesteps, while in the real world we collect 5 hours of data with asynchronous data collection and training on a single NVIDIA RTX 6000 Ada GPU. To match real-world conditions, episodes in simulation proceed continuously without manual resets to predetermined states. Additional training details and all hyperparameters are provided in Appendix D. Demonstration videos are available at adaptive-intelligent-robotics.github.io/URSA.
@@ 5.2 RQ1: Unsupervised Discovery of Diverse and High-Performing Locomotion Skills
Figure 3 illustrates the distribution of average joint angles, evaluated in simulation, across different leg combinations. URSA discovers a wide variety of motion patterns through varied use of the leg joints. This is further supported by Figure 1, which visualizes four example behaviors from the repertoire, ranging from gaits that keep the torso near the ground to more agile, coordinated walking motions. By contrast, DayDreamer converges on a single locomotion strategy that maximizes return, while DOMiNiC—though able to learn multiple skills, enforces near-optimality, resulting in less joint-level diversity across its 8 discrete behaviors. Quantitatively, when projecting the joint angle trajectories onto discretized spaces of average angles, URSA achieves 4 times greater coverage compared to the baselines, demonstrating significantly more behavioral diversity (Fig. 3).
@@ 5.3 RQ2: Downstream Use of Skills for Damage Adaptation
To evaluate utility of URSA’s learned repertoire for downstream tasks, we study its effectiveness in adapting to physical damage across simulation (Figure 4) and real-world (Figure 5) scenarios. We apply damage scenarios including single joint failures and full failures of one or two legs [^1] Joint notation: F = Front, B = Back, R = Right, L = Left (e.g., FR = Front Right, BL = Back Left)..
@@
In Figure 5, we evaluate only URSA ++ ITE due to the impracticality of executing all skills. URSA outperforms DayDreamer in all damage scenarios except for back-right leg failures (left panel). Notably, URSA demonstrates reduced overall performance degradation under damage, leveraging its diverse skill repertoire to compensate for broken joints. This is evident in Figure 5 (middle), where URSA ++ ITE progressively improves total return after consecutive iterations, recovering within 8 iterations during back-left leg damage. The underlying skill repertoire’s quality is shown in Figure 5 (right), where consistent improvement in average rewards indicates that URSA builds a diverse set of high-performing behaviors. In contrast, DayDreamer tends to overfit to forward locomotion strategies, leading to significant performance drops when the front legs are damaged. Additionally, URSA achieves higher return in the undamaged setting, suggesting that behavioral diversity not only improves robustness but also acts as a stepping stone toward discovering high-performing locomotion strategies, as opposed to single-skill optimization.
@@ 5.4 RQ3: Heuristic-Based Skill Discovery for Velocity Tracking
We assess the usefulness of the discovered skills by measuring how well they track target velocity commands. Figure 6 reports the tracking error across a grid of desired forward and angular velocities. URSA learns a repertoire that spans a broad range of these velocities, as confirmed by the plotted skill distribution (right). However, we observe that tracking performance degrades as target velocities deviate further from zero, particularly in the case of forward velocity. This discrepancy likely arises because some target velocities, while instantaneously achievable, cannot be consistently sustained across the entire duration of a skill. These findings highlight that URSA is capable not only of discovering diverse behaviors, but also of learning skill representations that are suitable for real-world, target-driven control.
@@ 6 Related Work
Recent methods aim to discover diverse behaviors without explicit rewards. Unsupervised RL approaches [23, 24, 25, 26] aim at maximizing the mutual information between abstract skills and trajectories to induce distinct behaviors, while SMERL [27] and DOMiNO [28] improve robustness by learning multiple solutions per task. In the Quality-Diversity (QD) domain, methods such as TAXONS [29], STAX [30], and AURORA [10] learn behavioral descriptors directly from data, removing the need for hand-crafted behavioral descriptors. However, these approaches heavily rely on simulation and face challenges in real-world applications. Our work builds on these methods by introducing safety constraints and efficient skill sampling for real-world learning, as well as imagination-based planning to reduce physical interactions.
@@
Real-world learning faces challenges like including limited data, safety concerns, and the lack of reset mechanisms. Reset-free QD [31, 7] selects behaviors that self-reset. Laversanne-Finot et al. [32] demonstrate diverse skill acquisition in robotic arms via goal-directed exploration, but still relies on episodic resets. off-DADS [6] enables skill discovery on real quadrupeds with off-policy RL, but does not consider any extrinsic reward. DayDreamer [13] sets a standard for sample-efficient learning using world models, and similarly, SERL [1], A Walk in the Park [2] and CrossQ [33] improve sample efficiency through careful system and algorithmic innovations. Our approach, like DayDreamer, leverages world models for unsupervised multi-skill discovery, adding safety constraints, and sophisticated skill selection mechanisms.
@@ Appendix A Supplementary Analysis
URSA does not consistently outperform all baselines across individual damage scenarios (Fig. 4). Nevertheless, URSA and URSA+ITE exhibit less performance degradation overall, indicating greater robustness. This is illustrated by the complementary cumulative distribution functions (CCDF) of the median performance across our 9 damage scenarios considered (Fig. 7).
@@ B.2 Kernel Density Estimation Sampling
Initially, the repertoire $\mathcal{R}$ is empty, so we sample skills from a standard normal distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. These initial random skills do not compromise learning since sampled skills are only used for action selection and policy updates, and are never stored in the dataset $\mathcal{D}$. Once the repertoire contains enough skills for a non-degenerate covariance matrix (usually when $|\mathcal{R}|\geq D$, where $D$ is the dimensionality of $\boldsymbol{\phi}$), we transition to sampling from the KDE fitted on $\mathcal{R}$.
@@
The covariance matrix $\boldsymbol{\Sigma}$} is the empirical covariance of all skills in the repertoire, scaled by Scott’s rule bandwidth $h=N_{\mathcal{R}}^{-\frac{1}{D+4}}$, such that the final covariance is ${\boldsymbol{\Sigma}}=h^{2}{\boldsymbol{\Sigma}}_{\text{empirical}}$. This sampling procedure naturally concentrates samples around regions of high skill density while allowing exploration of nearby skill variations through the Gaussian kernels.
@@ D.1 Unsupervised Quality-Diversity
This penalizes instability and unsafe motion. We aim to learn a repertoire of $N_{\boldsymbol{z}}=512$ distinct skills in this setting.
D.2 ITE for Damage Adaption
In this section, we provide the implementation details of the iterative trial and error (ITE) algorithm, adapted for use with URSA’s skill repertoire. The approach is based on the methodology introduced in Cully et al. (2015) [3].
@@ D.4 Hyperparameters
A full list of rollout parameters, and training constants used in our experiments is provided in Table 1.

### v3

@@ 1 Introduction
In this work, we introduce Unsupervised Real-world Skill Acquisition (URSA), a framework for unsupervised QD that operates directly on physical robots. URSA extends the Quality-Diversity Actor-Critic (QDAC) framework [13] for unsupervised learning and leverages the DayDreamer [14] world model as its backbone, enabling efficient skill acquisition directly on hardware without relying on simulation or sim-to-real transfer. We evaluate URSA on a Unitree A1 quadruped and show it outperforms baselines in 5 out of 9 simulated and 3 out of 5 real-world damage scenarios. This demonstrates its potential for enabling adaptable, resilient skill discovery directly on physical robots. Our key contributions are:
@@ 4.2 Efficient Skill Sampling from the Reachable Space
To maximize the diversity of achieved skills, URSA must efficiently explore the reachable skill space $\mathcal{Z}_{p}$ while ensuring safe execution. URSA addresses this by maintaining a diversity-aware repertoire $\mathcal{R}$ and sampling skills from a flexible surrogate distribution $q(\cdot)$ learned from experience. Figure 2 and Algorithm 1 provide high-level overviews of the skill collection and sampling process.
@@
As the repertoire expands, the typical distance between skills increases, requiring adaptive tuning of the constraint threshold $\delta$ in Problem P3. To that end, we introduce a hyperparameter $N_{\boldsymbol{z}}$, which controls the number of distinct skills that the robot can execute. We compute $\delta$ as the mean nearest-neighbor distance, if the repertoire contained $N_{\boldsymbol{z}}$ uniformly distributed skills (see equation in Appendix B.5).
This ensures that executing one skill does not inadvertently satisfy multiple constraints.
@@ 5.1 Experimental Setup and Baselines
We evaluate URSA on the Unitree A1 quadruped in both PyBullet simulation and real-world environments. The robot state includes joint angles and velocities for its 12 joints; actions specify target joint positions executed via a 20 Hz PD controller. Skills are sampled every 250 steps from a fixed repertoire of $N_{R}=4096$. Training in simulation runs for 1 million timesteps, while in the real world, we collect 5 hours of data with asynchronous data collection and training on a single NVIDIA RTX 6000 Ada GPU. To match real-world conditions, episodes in simulation proceed continuously without manual resets to predetermined states. Additional training details and all hyperparameters are provided in Appendix D. Demonstration videos are available at adaptive-intelligent-robotics.github.io/URSA.
@@ 5.2 RQ1: Unsupervised Discovery of Diverse and High-Performing Locomotion Skills
Figure 3 illustrates the distribution of average joint angles, evaluated in simulation, across different leg combinations. URSA discovers a wide variety of motion patterns through varied use of the leg joints. This is further supported by Figure 1, which visualizes four example behaviors from the repertoire, ranging from gaits that keep the torso near the ground to more agile, coordinated walking motions. By contrast, DayDreamer converges on a single locomotion strategy that maximizes return, while DOMiNiC—though able to learn multiple skills—enforces near-optimality, resulting in less joint-level diversity across its 8 discrete behaviors. Quantitatively, when projecting the joint angle trajectories onto discretized spaces of average angles, URSA achieves 4 times greater coverage compared to the baselines, demonstrating significantly more behavioral diversity (Fig. 3).
@@ 5.3 RQ2: Downstream Use of Skills for Damage Adaptation
To evaluate the utility of URSA’s learned repertoire for downstream tasks, we study its effectiveness in adapting to physical damage across simulation (Figure 4) and real-world (Figure 5) scenarios. We apply damage scenarios including single joint failures and full failures of one or two legs [^1] Joint notation: F = Front, B = Back, R = Right, L = Left (e.g., FR = Front Right, BL = Back Left)..
@@
In Figure 5, we evaluate only URSA ++ ITE due to the impracticality of executing all skills. URSA outperforms DayDreamer in all damage scenarios except for back-right leg failures (left panel). Notably, URSA demonstrates reduced overall performance degradation under damage, leveraging its diverse skill repertoire to compensate for broken joints. This is evident in Figure 5 (middle), where URSA ++ ITE progressively improves total return after consecutive iterations, recovering within 8 iterations during back-left leg damage. The underlying skill repertoire’s quality is shown in Figure 5 (right), where consistent improvement in average rewards indicates that URSA builds a diverse set of high-performing behaviors. In contrast, DayDreamer tends to overfit to forward locomotion strategies, leading to significant performance drops when the front legs are damaged. Additionally, URSA achieves higher returns in the undamaged setting, suggesting that behavioral diversity not only improves robustness but also acts as a stepping stone toward discovering high-performing locomotion strategies, as opposed to single-skill optimization.
@@ 5.4 RQ3: Heuristic-Based Skill Discovery for Velocity Tracking
We assess the usefulness of the discovered skills by measuring how well they track target velocity commands. Figure 6 reports the tracking error across a grid of desired forward and angular velocities. URSA learns a repertoire that spans a broad range of these velocities, as confirmed by the plotted skill distribution (Fig. 6, left). However, we observe that tracking performance degrades as target velocities deviate further from zero, particularly in the case of forward velocity. This discrepancy likely arises because some target velocities, while instantaneously achievable, cannot be consistently sustained across the entire duration of a skill. These findings highlight that URSA is capable not only of discovering diverse behaviors, but also of learning skill representations that are suitable for real-world, target-driven control.
@@ 6 Related Work
Recent methods aim to discover diverse behaviors without explicit rewards. Unsupervised RL approaches [23, 24, 25, 26] aim at maximizing the mutual information between abstract skills and trajectories to induce distinct behaviors, while SMERL [27] and DOMiNO [28] improve robustness by learning multiple solutions per task. In the Quality-Diversity (QD) domain, methods such as TAXONS [29], STAX [30], and AURORA [10] learn behavioral descriptors directly from data, removing the need for hand-crafted behavioral descriptors. However, these approaches heavily rely on simulation and face challenges in real-world applications. Our work builds on these methods by introducing safety constraints and efficient skill sampling for real-world learning, as well as imagination-based training to reduce physical interactions.
@@
Real-world learning faces challenges like including limited data, safety concerns, and the lack of reset mechanisms. Reset-free QD [31, 7] selects behaviors that self-reset. Laversanne-Finot et al. [32] demonstrate diverse skill acquisition in robotic arms via goal-directed exploration, but still rely on episodic resets. off-DADS [6] enables skill discovery on real quadrupeds with off-policy RL, but does not consider any extrinsic reward. DayDreamer [14] sets a standard for sample-efficient learning using world models, and similarly, SERL [1], A Walk in the Park [2] and CrossQ [33] improve sample efficiency through careful system and algorithmic innovations. Our approach, like DayDreamer, leverages world models for unsupervised multi-skill discovery, adding safety constraints, and sophisticated skill selection mechanisms.
@@ Appendix A Supplementary Analysis
URSA does not consistently outperform all baselines across individual damage scenarios (Fig. 4). Nevertheless, URSA and URSA ++ ITE exhibit less performance degradation overall, indicating greater robustness. This is illustrated by the complementary cumulative distribution functions (CCDF) of the median performance across our 9 damage scenarios considered (Fig. 7).
@@ B.2 Kernel Density Estimation Sampling
Initially, the repertoire $\mathcal{R}$ is empty, so we sample skills from a standard normal distribution $\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. These initial random skills do not compromise learning since sampled skills are only used for action selection and policy updates, and are never stored in the dataset $\mathcal{D}$. Once the repertoire contains enough skills for a non-degenerate covariance matrix, we transition to sampling from the KDE fitted on $\mathcal{R}$.
@@
The covariance matrix $\boldsymbol{\Sigma}$} is the empirical covariance of all skills in the repertoire scaled by Scott’s rule bandwidth $h=N_{\mathcal{R}}^{-\frac{1}{D+4}}$, such that the final covariance is ${\boldsymbol{\Sigma}}=h^{2}{\boldsymbol{\Sigma}}_{\text{empirical}}$. This sampling procedure naturally concentrates samples around regions of high skill density while allowing exploration of nearby skill variations through the Gaussian kernels.
@@ D.1 Unsupervised Quality-Diversity
This penalizes instability and unsafe motion when the robot deviates from an upright posture. We aim to learn a repertoire of $N_{\boldsymbol{z}}=512$ distinct skills in this setting.
D.2 ITE for Damage Adaptation
In this section, we provide the implementation details of the iterative trial and error (ITE) algorithm, adapted for use with URSA’s skill repertoire. The approach is based on the methodology introduced in Cully et al. [3].
@@ D.4 Hyperparameters
A full list of rollout parameters and training constants used in our experiments is provided in Table 1.